# MalMixer: Few-Shot Malware Classification with Retrieval-Augmented Semi-Supervised Learning

### v2

@@ IV Motivating Example
The three assumptions described in Section III-C underpin the logic behind several key steps in our augmentation pipeline. Next, we provide a case study in which each step of our augmentation pipeline is discussed. In doing so, we aim to intuitively explain how each step works together to form novel, domain-knowledge-aware synthetic malware samples. Such augmented synthetic samples can then effectively enhance the performance of the overarching semi-supervised framework.
@@
We first consider a pair of similar real malware Samples A and B, where sample A is labeled, but B is not. Both samples belong to the same family, unruy, a family of trojan adware samples. Conventional augmentation approaches that mix or perturb samples only consider feature values without regard for the underlying semantics of those features. For example, each of Sample A and B contain a MurmurHash3 hash of the loaded libraries. Naive approaches mutate these samples by averaging or adding noise to these hash values, resulting in values that do not correspond to sensible input strings.
@@
We also assess the compatibility between interpolatable and non-interpolatable features of the synthetic sample, providing intuition for our alignment method. A substantial number of key interpolatable features of the synthetic sample (e.g., the number of import functions, import table size, import address table size, and MEM_WRITE, MEM_READ, and MEM_EXECUTE operations) match exactly those of the real sample from which we retrieved its non-interpolatable features. Other features, such as Shannon’s entropy of printable characters and the byte histogram, differ by only around $10\%$ from the real sample. This suggests that our augmentation method creates a synthetic sample that contains realistic values for a novel and plausible unruy executable.
In addition to this intuitive example, we also conduct an Ablation Study in Section VII-E to empirically showcase how each step in our proposed augmentation pipeline contributes to accurate few-shot malware classification.
V Experimental Settings
In this section, we present an overview of the datasets we use to evaluate MalMixer. We also include our model architectures, implementation, training hyperparameters, and hardware resources in Section B in the Appendix.
V-A Training Datasets
@@
The BODMAS dataset contains $57,293$ malware samples representing $581$ families, obtained and timestamped between August 2019 and September 2020. The MOTIF dataset contains $3,095$ disarmed PE malware samples from $454$ families, published between January 2016 and January 2021. All samples in both datasets are represented by a 2381-dimensional feature representation statically extracted from Windows PE binaries. The static features encompass byte entropy histogram [50], string entropy, PE header information, sections table information, and hashed imported libraries, among others. The set of features in BODMAS and MOTIF (thus used in MalMixer) is the same as EMBER, and feature extraction is conducted through the LIEF project [51], a lightweight PE parser. In general, the static features used are not engineered for detecting certain families but rather provide a comprehensive summary of a PE file. We elided certain feature extraction and preprocessing steps as they can be found in detail in the EMBER dataset [42]. Below, we discuss preprocessing steps and subsets of BODMAS and MOTIF we use for particular experiments.
@@
V-B Motivation for dataset pre-processing
We train and evaluate MalMixer on subsets of BODMAS and MOTIF consisting of all samples in the top 20 to top 50 most prevalent malware families. We justify these choices as follows: First, as shown in Table I, MalMixer can perform multi-way classification with a large number of classes. We further provide two additional experiments in Section D in the Appendix to demonstrate MalMixer’s effectiveness on the entire BODMAS dataset. In BODMAS, the top 20 to 30 most prevalent families account for $69.42\%$ to $80.02\%$ of all malware samples. In MOTIF, the top 50 most prevalent families account for $50.50\%$ of all samples. Therefore, selecting these top families strikes a balance between covering a diverse subset of BODMAS/MOTIF and ensuring experimental expediency. Second, most existing ML-based classifiers are typically evaluated on datasets with fewer than 25 classes [30, 22, 52, 53, 54]. By conducting 20-to-50-way classification, we can sufficiently demonstrate the effectiveness of our model in settings that follow current best practices. Third, MalMixer’s model design assumes that practitioners have pre-identified families of interest and provided at least one labeled sample from each family. In practice, if a novel samples from an unseen family appear in the wild, MalMixer requires only a small number of labeled examples of that new family to retrain the model, as demonstrated in Section VII-D. We devise several datasets based on BODMAS and MOTIF to evaluate our classifier, each described below.
TABLE II: BODMAS-20, MOTIF-50, & BODMAS-Temporal statistics.
@@ 30 Zero-shot 36677 9170 10 4090
One-shot - - 0 0
BODMAS-20 and MOTIF-50: We curated BODMAS-20 and MOTIF-50 to evaluate the performance of MalMixer in standardized semi-supervised classification settings. BODMAS-20 consists of all samples within the 20 most prevalent families in BODMAS, and MOTIF-50 consists of all samples within the 50 most prevalent families in MOTIF. BODMAS-20 is relatively balanced with the largest family, sfone, consisting of 4729 samples and the smallest family, unruy, consists of 865 samples. Within MOTIF-50, the family sizes range from 142 to 16 samples. BODMAS-20 and MOTIF-50 both use an $80\%/20\%$ train/test split.
BODMAS-Temporal: We curate BODMAS-Temporal to evaluate the performance of MalMixer when all testing families are represented in the training set. However, the training and testing splits are taken from non-overlapping points in time (based upon dates in BODMAS), thus potentially capturing concept drift where distributions shift over time [25]. All samples from the top 20 most prevalent families are included. However, the training set only consists of samples dated before 2020. Most of the test set consists of samples gathered at a later time from a distinct distribution [25]. Specifically, five test sets are created: timestampped before 2020 (for a baseline), each quarter during 2020 (Q1, Q2, and Q3), and all samples from 2020. Each train and test set contains at least one sample from each of the 20 families. This dataset uses a 80%/20% train/test split. Table II summarizes the statistics of BODMAS-20 and BODMAS-Temporal.
BODMAS-LeaveOut: BODMAS-LeaveOut is curated to evaluate MalMixer in handling malware from previously unlabeled families. To this end, we consider three training set-ups: (1) the Top 21 most prevalent families are selected with one family deemed ‘unidentified’ during training; (2) Top 25 families with five ‘unidentified’; and (3) Top 30 families with ten ‘unidentified’. As illustrated by Table III, each of these three testing set-ups contain two scenarios: ‘Zero-shot’ and ‘One-shot’. In ‘Zero-shot,’ families deemed ‘unidentified’ are not labeled in training. In ‘One-shot,’ we introduce a single labeled instance for each originally ‘unidentified’ family during training, so that all families become newly identified. The ‘unidentified’ families and the one-shot examples are selected randomly. The train-test split is performed before the ‘unidentified’ families are selected, so that samples from all families are present in both train and test sets. Note that by choosing these set-ups, it remains constant that 20 families are always ‘identified,’ and thus labeled in training. This design enables the assessment of MalMixer when faced with completely unseen malware families, as well as its capacity to be retrained to handle previously-unlabeled families for which only one sample is newly-labeled.
VI Evaluation Design
@@
    RQ1: How well does MalMixer sustain satisfactory performance with decreasing labeled training data in semi-supervised malware family classification?
@@
    RQ4: How effective is MalMixer when classifying samples from families unidentified during training?
@@
In this section, we discuss the methodology for answering each RQ. Then, Section VII presents the empirical results.
@@
RQ1: Saturation Analysis First, we assess how well MalMixer can maintain satisfactory performance with decreasing amounts of labeled data through saturation analysis using BODMAS-20. This analysis directly addresses a key challenge in classifying evolving malware: the trade-off between the time and human resources needed to create larger labeled datasets vs. immediate, satisfactory model accuracy.
@@
We provide comprehensive evaluation of this trade-off by randomly selecting $0.1\%$, $0.5\%$, $1\%$, $2\%$, $5\%$, $10\%$, $50\%$, and $75\%$ of all data in training set as labeled data, ensuring that at least one sample from each family is labeled. In the $0.1\%$ setting, no family contains more than 5 labeled samples.
@@
RQ2: Model Comparison MalMixer aims to enhance malware classification using static malware features, an area of significant research importance [3, 43]. Thus, we measure MalMixer’s relative performance in standardized few-shot settings by comparing with existing static-feature-based semi-supervised learning models on BODMAS-20 and MOTIF-50. We consider the performance of six baseline semi-supervised methods: Pseudo-Label [55], Virtual Adversarial Training (VAT) [56], Virtual Adversarial Training equipped with our proposed retrieval-augmented malware data (VAT aug), MixUp [37], MixMatch without augmentation (MixMatch) [24], and MixMatch applied with Gaussian-noise-injection augmentation (MixMatch Gaussian). These techniques encompass a swath of state-of-the-art techniques as baselines for comparison. Specifically, Pseudo-Label uses entropy-minimization by iteratively creating hard labels on high-confidence predictions, which are then incorporated into the labeled set. In contrast, VAT strategically perturbs the labeled data to maximize changes in output class distributions, making the model more robust against variations from the training set. Although MixUp is traditionally considered a regularization method for supervised learning [24], we adapt the model and empirically verify that it is well-suited for few-shot malware classification tasks and thus incorporate it as a baseline method. Last, MixMatch is a widely-used image-based semi-supervised learning method. Its overarching framework aligns closely with MalMixer, though MalMixer’s augmentation approach innovates through more sensible feature manipulations via domain-knowledge-awareness. Given that the traditional image-based augmentation techniques are not directly applicable to malware, we adopt two variants of MixMatch: one without any augmentation and another with Gaussian noise injection. In the latter, the noise is random, drawing from the Gaussian distribution of each family in the labeled set. This set-up allows for a fair comparison with MalMixer, which introduces novel augmentation strategies tailored for malware classification. When evaluating on BODMAS-20, we consider $0.1\%$, $0.5\%$, $1\%$, $2\%$ of training set as labeled data. On MOTIF-50, we consider $1\%$ of the training set as labeled. We ensure at least one sample from each family is labeled.
@@
RQ3: Temporal Analysis Our third task is to evaluate MalMixer’s robustness against temporal shift of malware distribution through leveraging test samples collected from different periods of time than the training samples. We use BODMAS-Temporal, selecting $1\%$ of the training set as labeled data because it yields a desired trade-off between the number of labeled samples and model accuracy (discussed in Section VII-A). We compare MalMixer’s performance with the six other baseline methods mentioned above.
RQ4: Leave-Out Testing Next, we evaluate MalMixer’s ability to be quickly retrained to handle test samples representing malware families not originally identified in the training data. Our experiment simulates the arrival of a new batch of unlabeled malware containing mostly previously identified families but at least one unidentified family. Now, we perform a random train/test split on the whole batch (agnostic of family). After the train/test split, samples from all families are present in both train and test sets. However, at this point, the unidentified family in the training set remains completely unlabeled and cannot be used as helpful training data. We aim to show that MalMixer can address this issue by accurately classifying the previously-unidentified family as long as one sample within that family becomes provided, reducing the amount of labeling required to identify a new family.
To simulate this, we employ BODMAS-LeaveOut. For each family deemed ‘identified’ in training, we select $1\%$ of the training set as labeled data. We first evaluate MalMixer in the ‘Zero-shot’ scenario, where the ‘unidentified’ families are not labeled in training. Then, we retrain our model in the ‘One-shot’ scenario, introducing a single labeled instance for each previously ‘unidentified’ family. This retraining allows us to quantify MalMixer’s ability to classify newly-labeled malware samples that belong to previously unlabeled families.
VII Results
@@
In this section, we present an evaluation of MalMixer. We discuss each research question through: (RQ1) a saturation analysis to demonstrate model performance with respect to the volume of labeled malware samples; (RQ2) a model comparison with state-of-the-art models and technqiues; (RQ3) an evaluation of model robustness when classifying new samples from a later period of time than samples used in training; and (RQ4) an evaluation of model robustness when classifying samples in families that were unlabeled during training.
@@
TABLE IV: Saturation evaluation of MalMixer against BODMAS-20, with ResNet-12 provided as a supervised standard for reference.
@@
ResNet-12 [57] $100\%$ 93.45 93.72 94.07 93.87
TABLE V: Comparison of MalMixer against semi-supervised methods on BODMAS-20 ($0.1\%$ labeled data) and MOTIF-50 ($1\%$ labeled data).

### v4

@@ 4 Motivating Example
The three assumptions described in Section 3.3 underpin the logic behind several key steps in our data augmentation pipeline. Next, we provide a case study in which each step of our data augmentation pipeline is discussed. In doing so, we aim to intuitively explain how each step works together to form novel, domain-knowledge-aware synthetic malware samples. Such augmented synthetic samples can then effectively enhance the performance of the overarching semi-supervised framework.
@@
We first consider a pair of similar real malware Samples A and B, where sample A is labeled, but B is not. Both samples belong to the same family, unruy, a family of trojan adware samples. Conventional data augmentation approaches that mix or perturb samples only consider feature values without regard for the underlying semantics of those features. For example, each of Sample A and B contains a MurmurHash3 hash of the loaded libraries. Naive approaches mutate these samples by averaging or adding noise to these hash values, resulting in values that do not correspond to sensible input strings.
@@
We also assess the compatibility between interpolatable and non-interpolatable features of the synthetic sample, providing intuition for our alignment method. A substantial number of key interpolatable features of the synthetic sample (e.g., the number of import functions, import table size, import address table size, and MEM_WRITE, MEM_READ, and MEM_EXECUTE operations) match exactly those of the real sample from which we retrieved its non-interpolatable features. Other features, such as Shannon’s entropy of printable characters and the byte histogram, differ by only around $10\%$ from the real sample. This suggests that our data augmentation method creates a synthetic sample that contains realistic values for a novel and plausible unruy executable.
In addition to this intuitive example, we also conduct an Ablation Study in Section 7.5 to empirically showcase how each step in our proposed augmentation pipeline contributes to accurate few-shot malware classification.
5 Experimental Settings
In this section, we present an overview of the datasets we use to evaluate MalMixer. We also include our model architectures, implementation, training hyperparameters, and hardware resources in Appendix A.
5.1 Training Datasets
@@
The BODMAS dataset contains $57,293$ malware samples representing $581$ families, obtained and timestamped between August 2019 and September 2020. The MOTIF dataset contains $3,095$ disarmed PE malware samples from $454$ families, published between January 2016 and January 2021. All samples in both datasets are represented by a 2381-dimensional feature representation statically extracted from Windows PE binaries. The static features encompass byte entropy histogram [50], string entropy, PE header information, sections table information, and hashed imported libraries, among others. The set of features in BODMAS and MOTIF (thus used in MalMixer) is the same as EMBER, and feature extraction is conducted through the LIEF project [51], a lightweight PE parser. In general, the static features used are not engineered for detecting certain families but rather provide a comprehensive summary of a PE file. We elided certain feature extraction and pre-processing steps as they can be found in detail in the EMBER dataset [42]. Below, we discuss pre-processing steps and subsets of BODMAS and MOTIF we use for particular experiments.
@@
5.2 Motivation for dataset pre-processing
We train and evaluate MalMixer on subsets of BODMAS and MOTIF consisting of all samples in the top 20 to top 50 most prevalent malware families. We justify these choices as follows: First, as shown in Table I, MalMixer can perform multi-way classification with a large number of classes. We further provide two additional experiments in Section E in the Appendix to demonstrate MalMixer’s effectiveness on the entire BODMAS dataset. In BODMAS, the top 20 to 30 most prevalent families account for $69.42\%$ to $80.02\%$ of all malware samples. In MOTIF, the top 50 most prevalent families account for $50.50\%$ of all samples. Therefore, selecting these top families strikes a balance between covering a diverse subset of BODMAS/MOTIF and ensuring experimental expediency. Second, most existing ML-based classifiers are typically evaluated on datasets with fewer than 25 classes [29, 22, 52, 53, 54]. By conducting 20-to-50-way classification, we can sufficiently demonstrate the effectiveness of our model in settings that follow current best practices. Third, MalMixer’s model design assumes that practitioners have pre-identified families of interest and provided at least one labeled sample from each family. In practice, if novel samples from an unseen family appear in the wild, MalMixer requires only a small number of labeled examples of that new family to retrain the model, as demonstrated in Section 7.4. We devise several datasets based on BODMAS and MOTIF to evaluate our classifier, each described below.
TABLE II: Dataset Statistics of BODMAS-20, MOTIF-50, & BODMAS-Temporal.
@@ 30 Zero-shot 36677 9170 10 4090
One-shot 36677 9170 0 0
BODMAS-20 and MOTIF-50: We curated BODMAS-20 and MOTIF-50 to evaluate the performance of MalMixer in standardized semi-supervised classification settings. BODMAS-20 consists of all samples within the 20 most prevalent families in BODMAS, and MOTIF-50 consists of all samples within the 50 most prevalent families in MOTIF. BODMAS-20 is relatively balanced with the largest family, sfone, consisting of 4729 samples and the smallest family, unruy, consists of 865 samples. Within MOTIF-50, the family sizes range from 142 to 16 samples. BODMAS-20 and MOTIF-50 both use an $80\%/20\%$ train/test split. Additionally, since MalMixer relies on retrieving non-interpolatable features, Appendix H also provides dataset statistics on train-test overlap of non-interpolatable features.
BODMAS-Temporal: We curate BODMAS-Temporal to evaluate the performance of MalMixer when all testing families are represented in the training set. However, the training and testing splits are taken from non-overlapping points in time (based upon dates in BODMAS), thus potentially capturing concept drift where distributions shift over time [24]. All samples from the top 20 most prevalent families are included. However, the training set only consists of samples dated before 2020. Most of the test set consists of samples gathered at a later time from a distinct distribution [24]. Specifically, five test sets are created: timestamped before 2020 (for a baseline), each quarter during 2020 (Q1, Q2, and Q3), and all samples from 2020. Each train and test set contains at least one sample from each of the 20 families. This dataset uses a 80%/20% train/test split on samples dated before 2020. Table II summarizes the statistics of BODMAS-20, MOTIF-50, and BODMAS-Temporal.
BODMAS-LeaveOut: BODMAS-LeaveOut is curated to evaluate MalMixer in handling malware from previously unlabeled families. To this end, we consider three training set-ups: (1) the Top 21 most prevalent families are selected with one family deemed ‘unidentified’ during training; (2) Top 25 families with five ‘unidentified’; and (3) Top 30 families with ten ‘unidentified’. As illustrated by Table III, each of these three testing set-ups contains two scenarios: ‘Zero-shot’ and ‘One-shot’. In ‘Zero-shot,’ families deemed ‘unidentified’ are not labeled in training. In ‘One-shot,’ we introduce a single labeled instance for each originally ‘unidentified’ family during training, so that all families become newly identified. The ‘unidentified’ families and the one-shot examples are selected randomly. The train-test split is performed before the ‘unidentified’ families are selected, so that samples from all families are present in both train and test sets. Note that by choosing these set-ups, it remains constant that 20 families are always ‘identified,’ and thus labeled in training. This design enables the assessment of MalMixer when faced with completely unseen malware families, as well as its capacity to be retrained to handle previously-unlabeled families for which only one sample is newly-labeled.
6 Evaluation Design
@@
    RQ1: How well does MalMixer sustain satisfactory performance with decreasing labeled training data in semi-supervised malware classification?
@@
    RQ4: How effective is MalMixer when classifying samples from families unidentified in training?
@@
In this section, we discuss the methodology for answering each RQ. Section 7 presents the empirical results.
@@
RQ1: Saturation Analysis First, we assess how well MalMixer can maintain satisfactory performance with decreasing amounts of labeled data through saturation analysis. This analysis directly addresses a key challenge in classifying evolving malware: the trade-off between the time and human resources needed to create larger labeled datasets vs. immediate, satisfactory model accuracy.
@@
Using BODMAS-20, we provide comprehensive evaluation of this trade-off by randomly selecting $0.1\%$, $0.5\%$, $1\%$, $2\%$, $5\%$, $10\%$, $50\%$, and $75\%$ of all data in training set as labeled data, ensuring that at least one sample from each family is labeled. In the $0.1\%$ setting, no family contains more than 5 labeled samples.
@@
RQ2: Model Comparison MalMixer aims to enhance malware classification using static malware features, an area of significant research importance [3, 43]. Thus, we measure MalMixer’s relative performance in standardized few-shot settings by comparing with existing static-feature-based semi-supervised learning models on BODMAS-20 and MOTIF-50. We consider the performance of seven baseline semi-supervised methods: Pseudo-Label [55], Virtual Adversarial Training (VAT) [56], Virtual Adversarial Training equipped with our proposed retrieval-augmented pseudo-malware data (VAT aug), MixUp [36], MixMatch without data augmentation (MixMatch) [23], MixMatch applied with Gaussian-noise-injection data augmentation (MixMatch Gaussian), and MORSE [47]. These techniques encompass a swath of state-of-the-art techniques as baselines for comparison. Specifically, Pseudo-Label uses entropy-minimization by iteratively creating hard labels on high-confidence predictions, which are then incorporated into the labeled set. In contrast, VAT strategically perturbs the labeled data to maximize changes in output class distributions, making the model more robust against variations from the training set. Although MixUp is traditionally considered a regularization method for supervised learning [23], we adapt the model and empirically verify that it is well-suited for few-shot malware classification tasks and thus incorporate it as a baseline method. Last, MixMatch is a widely-used image-based semi-supervised learning method. Its overarching framework aligns closely with MalMixer, though MalMixer’s augmentation approach innovates through more sensible feature manipulations via domain-knowledge-awareness. Given that the traditional image-based augmentation techniques are not directly applicable to malware, we adopt two variants of MixMatch: one without any augmentation and another with Gaussian noise injection. In the latter, the noise is random, drawing from the Gaussian distribution of each family in the labeled set. This set-up allows for a fair comparison with MalMixer, which introduces novel augmentation strategies tailored for malware classification. Finally, MORSE is a state-of-the-art malware family classification model originally designed to mitigate label noise and class imbalance through semi-supervised learning. Given its demonstrated effectiveness in label-scarce scenarios, its semi-supervised architecture, and its malware-specific data augmentation techniques, MORSE can be adapted to our few-shot setting with appropriate modifications. Refer to Appendix I for more details on MORSE and its adaptations.
When evaluating on BODMAS-20, we consider $0.1\%$ of training set as labeled data. On MOTIF-50, we consider $1\%$ of the training set as labeled. We ensure at least one sample from each family is labeled, and highlight that these settings are close to the one-shot regime.
@@
RQ3: Temporal Analysis Our third task is to evaluate MalMixer’s robustness against temporal shift of malware distribution through leveraging test samples collected from different periods of time than the training samples. We use BODMAS-Temporal, selecting $1\%$ of the training set as labeled data because it yields a desired trade-off between the number of labeled samples and model accuracy (discussed in Section 7.1). We compare MalMixer’s performance with the seven other baseline methods mentioned above.
RQ4: Leave-Out Testing Next, we evaluate MalMixer’s ability to quickly retrain to handle test samples representing malware families not originally identified in training. Like many ML malware classifiers [47, 24, 46], MalMixer relies on a closed-world assumption, whereby it can only classify samples into families labeled in training. In real-world settings, when a new malware batch arrives containing at least one previously unseen family, practitioners using closed-world models can first leverage existing methods (e.g. [57]) to reject unfamiliar samples and flag them for manual labeling. Once new families are labeled, we can retrain the model under a closed-world framework.
MalMixer’s few-shot ability allows it to be promptly retrained to classify new families as soon as one sample from each previously unidentified family is labeled. To simulate how effective MalMixer handles this retraining, we employ BODMAS-LeaveOut, which features a mix of predominantly ‘identified’ families alongside at least one ‘unidentified’ family. For each family deemed ‘identified’ in training, we randomly select $1\%$ of the training set as labeled data. In the ‘Zero-shot’ settings, we perform a random train/test split on the whole batch, after which samples from all families are present in both train and test sets. However, at this point, the ‘unidentified’ families in the training set remain completely unlabeled and cannot be used as helpful training data. This setting provides a baseline for how MalMixer (without retraining) would have performed under the existence of unseen families, potentially misclassifying them into existing families. In contrast, the ‘One-shot’ settings simulate the addition of a single labeled instance for each previously ‘unidentified’ family, after which MalMixer is retrained. The performance improvement observed between the ‘Zero-shot’ and ‘One-shot’ scenarios quantifies MalMixer’s ability to correctly classify newly labeled malware samples with minimal labeling effort.
7 Results
@@
In this section, we present an evaluation of MalMixer. We discuss each research question through: (RQ1) a saturation analysis to demonstrate model performance with respect to the volume of labeled malware samples; (RQ2) a model comparison with state-of-the-art models and techniques; (RQ3) an evaluation of model robustness when classifying new samples from a later period of time than samples used in training; and (RQ4) an evaluation of model retraining when classifying samples in families that were previously unlabeled.
@@
TABLE IV: Saturation evaluation of MalMixer against BODMAS-20, with Fully Connected ResNet and LightGBM provided as a supervised standard for reference.
@@
FC-ResNet [58] $100\%$ 93.45 93.72 94.07 93.87
LightGBM [59] $100\%$ 93.90 94.09 94.01 94.04
TABLE V: Comparison of MalMixer against semi-supervised methods on BODMAS-20 ($0.1\%$ labeled data) and MOTIF-50 ($1\%$ labeled data), where most families are identified by one-shot labels.